In a Newton's ring experiment, the diameter of the 10th ring changes from 1.40 cm to 1.27 cm when a liquid is introduced between the lens and the plate. To calculate the refractive index of the liquid.
1
Expert's answer
2019-11-26T10:45:11-0500

The radii of Newton's light rings with a constant radius of curvature of the lens can be found using the formula:

"r_k=\sqrt {\frac{(k-0.5)\cdot λ\cdot R}{n}}"

if 

"r_1 = 1.4 cm, r_2 = 1.27 cm"

"r_1=\sqrt {\frac{(10-0.5)\cdot λ\cdot R}{n_1}}, n_1=1"

"r_2=\sqrt {\frac{(10-0.5)\cdot λ\cdot R}{n_2}}"

Next, divide r1 by r2 and get:

"\frac{r_1}{r_2}=\sqrt{n_2}"

"n_2=\frac{r_{1}^2}{r_{2}^2}=\frac{1.4^2}{1.27^2}=1.22"